Stat 10 - (00770) - Fall 2018-Chaffey Homework: Hmwrk 12 Score: 0 of 1 pt 11.1.19-T n 1942, an organization surveyed 1100 adults and asked, Are you a total abstainer from, or do you on occasion consume, alcoholic beverages? Of the 1100 aduls juan saucedo 읊 : 12/19/18 1:09 AM 30f 4 complete) ▼ Hw Score: 50%, 2 of 4 pts abstainers. parts (a) and (b) below The proportions of the adults who took the 1942 survey and the recent survey who were total abstainers are 37 and 3, respectively. Round to three decimal places as needed.) (b) Has the proportion of adults who totally abstain from alcohol changed? Use the First verify the model requirements. Select all that apply. #0 05 level of significance JA np (1-1) 10 and n2pa (1-2) 210 10 and np(1-2) 2 10 B. The sample size is more than 5% of the population size for each sample. C. The samples are dependent D. The samples are independent. E. The sample size is less than 5% of the population size for each sample. F. The data come from a population that is normally distributed. Click to select your answer(s) and then click Check Answer 5 Clear All parts remaining
Answer #1

a)
1942

p^=X/n = 407/1100 = 0.37

recent survey
p^= 330/1100 = 0.3

b)
option A) , D) ,E) are correct